March 1, 201, Dorchester Company's beginning work in process inventory had 8,000 units. This is its only production department. Beginning WIP units were 50% complete as to conversion costs. Dorchester introduces direct materials at the beginning of the production process.  During March, all beginning WIP was completed and an additional 19,500 units were started and completed. Dorchester also started but did not complete 7,500 units. These units remained in ending WIP inventory and were 70% complete as to conversion costs. Dorchester uses the weighted average method. Use this information to determine for March 2019 the equivalent units of production for conversion costs. (Round & enter final answers to: the nearest whole dollar for total dollar answers, nearest penny for unit costs or nearest whole number for units)
Answer #1

Units completed and transferred out = Units in beginning WIP+Unit started and completed = 8000+19500 = 27500

Equivalent unit of conversion = Units completed and transferred out+(Ending WIP*percent completion)

= 27500+(7500*70%)

Equivalent unit of conversion = 32750 Units